# Automating the Diagnosis of Human Vision Disorders by Cross-modal 3D Generation

### v1

@@ 4.4 Two-stage 3D generation
To effectively leverage the projected high-level embeddings, we devise a two-stage 3D generation architecture characterized by progressively integrating high-level information during 3D object modeling. The first stage is a perceptual stage, utilizing the Neural Radiance Fields (NeRF) [34] for 3D representation. It defines an implicit 3D representation framework expressed as:
@@
This framework allows the gradients from the Score Distillation Sampling (SDS) [33] to directly influence the output $C(\mathbf{r})$, enabling simultaneous optimization of the density Multilayer Perceptron (MLP) $\sigma(\mathbf{r}(t))$ and color MLP $(\mathbf{r}(t),\mathbf{d})$. As a result, NeRF is particularly effective at modifying both the color and shape of the 3D scene in this stage. In our implementation, we incorporate Zero-1-to-3 [67] for low-level guidance and Stable Diffusion [31] for high-level guidance.
Subsequently, a semantic stage follows where we employ DMTet [68] to convert the above NeRF into a 3D mesh [69]. This conversion prepares the 3D model for further processing. We continue to harness Stable Diffusion [31], coupled with a depth-conditioned Control-Net [70] for high-level guidance. While the 3D mesh does not possess the same versatility as NeRF in shaping modifications, it excels in capturing detailed RGB color information. Therefore, after establishing the coarse shape in the perceptual stage, we shift our focus to refining the color details in this stage. This two-tier design allows for an intricate and nuanced development of the 3D scene, leading to a final output that is both elaborately detailed and of high resolution.
@@ 4.5 Training and inference
During training, we optimize the parameters for both low-level encoder and high-level encoder, leaving all the rest components frozen. We use a training dataset with image-text pairs. Inspired by MindEye [35], the loss function for low-level $L_{image}$ and high-level $L_{text}$ embeddings are designed with integration of bidirectional CLIP [72] contrastive learning and MixUp [71] augmentation to improve generalization in the beginning of training stage, as described in the follows:
@@
To analyze the generation quality, we introduce four fMRI-to-3D evaluation metrics: 3D Inception [36], 3D CLIP [37], 3D EfficientNet [38], and 3D Swav [39]. These metrics are designed to overcome the limitations of traditional pixel-by-pixel image reconstruction metrics like PixelCorrelation [76], which lacks robustness against rotations and translations in 3D space. For comprehensive evaluation, we construct a level-2 icosahedron with a 2.2 radius around the generated 3D scene and capture the 3D scene from all the icosahedron vertices to achieve multi-view coverage, similar to the approach in the text-to-3D benchmark [77]. The metrics 3D Inception and 3D CLIP, utilizing the final pooling layer of InceptionV3 [36] and CLIP model [37] respectively, assess the accuracy of two-way identification between original image embeddings and corresponding brain embeddings. Consequently, larger 3D Inception and 3D CLIP represent higher matching accuracy in features between ground truth images and generated 3D objects, indicating a well extracted features from fMRI. On the other hand, 3D EfficientNet and 3D Swav employ EfficientNet-B1 [38] and SwAV-ResNet50 [39] to calculate the average correlation distance between the extracted features. Lower correlation distance represents higher correlation between ground truth images and generated 3D objects, representing better reconstruction. In order to address view inconsistency in the final quality evaluation, we implement a regional convolution mechanism to smooth the scores across the local region, which treats the icosahedron as a graph composed of vertices and edges and performs mean pooling on this graph, employing a recursive formula:
@@ 6 Code Availability
Our github repository is https://github.com/fudan-zvg/Brain3D. Supplementary materials with video visualizations of more 3D generations are available at https://brain-3d.github.io/.

### v2

@@ 4.4 Two-stage 3D generation
To effectively leverage the projected high-level embeddings, we devise a two-stage 3D generation architecture characterized by progressively integrating high-level information during 3D object modeling(Fig. 6e). The first stage is a perceptual stage, utilizing the Neural Radiance Fields (NeRF) [46] for 3D representation. It defines an implicit 3D representation framework expressed as:
@@
This framework allows the gradients from the Score Distillation Sampling (SDS) [45] to directly influence the output $C(\mathbf{r})$, enabling simultaneous optimization of the density Multilayer Perceptron (MLP) $\sigma(\mathbf{r}(t))$ and color MLP $(\mathbf{r}(t),\mathbf{d})$. As a result, NeRF is particularly effective at modifying both the color and shape of the 3D object in this stage. In our implementation, we incorporate Zero-1-to-3 [79] for low-level guidance and Stable Diffusion [43] for high-level guidance.
Subsequently, a semantic stage follows where we employ DMTet [82] to convert the above NeRF into a 3D mesh [83]. This conversion prepares the 3D model for further processing. We continue to harness Stable Diffusion [43], coupled with a depth-conditioned Control-Net [84] for high-level guidance. While the 3D mesh does not possess the same versatility as NeRF in shaping modifications, it excels in capturing detailed RGB color information. Therefore, after establishing the coarse shape in the perceptual stage, we shift our focus to refining the color details in this stage. This two-tier design allows for an intricate and nuanced development of the 3D object, leading to a final output that is both elaborately detailed and of high resolution.
@@ 4.5 Training and inference
During training, we optimize the parameters for both low-level encoder and high-level encoder, leaving all the rest components frozen. The high-level encoder is detailed in Sec. 4.3, while the low-level encoder comprises MLP backbone, diffusion prior [80], and Versatile diffusion [81]. We use a training dataset with image-text pairs. Inspired by MindEye [47], the loss function for low-level $L_{image}$ and high-level $L_{text}$ embeddings are designed with integration of bidirectional CLIP [86] contrastive learning and MixUp [85] augmentation to improve generalization in the beginning of training stage, as described in the follows:
@@
To analyze the generation quality, we introduce four fMRI-to-3D evaluation metrics: 3D Inception [48], 3D CLIP [49], 3D EfficientNet [50], and 3D Swav [51]. These metrics are designed to overcome the limitations of traditional pixel-by-pixel image reconstruction metrics like PixelCorrelation [90], which lacks robustness against rotations and translations in 3D space. For comprehensive evaluation, we construct a level-2 icosahedron with a 2.2 radius around the generated 3D object and capture the 3D object from all the icosahedron vertices to achieve multi-view coverage, similar to the approach in the text-to-3D benchmark [91]. The metrics 3D Inception and 3D CLIP, utilizing the final pooling layer of InceptionV3 [48] and CLIP model [49] respectively, assess the accuracy of two-way identification between original image embeddings and corresponding brain embeddings. Consequently, larger 3D Inception and 3D CLIP represent higher matching accuracy in features between ground truth images and generated 3D objects, indicating a well extracted features from fMRI. On the other hand, 3D EfficientNet and 3D Swav employ EfficientNet-B1 [50] and SwAV-ResNet50 [51] to calculate the average correlation distance between the extracted features. Lower correlation distance represents higher correlation between ground truth images and generated 3D objects, representing better reconstruction. In order to address view inconsistency in the final quality evaluation, we implement a regional convolution mechanism to smooth the scores across the local region, which treats the icosahedron as a graph composed of vertices and edges and performs mean pooling on this graph, employing a recursive formula:
@@ 6 Code Availability
We enclosed our code and training steps in submission. Our github repository is https://github.com/fudan-zvg/Brain3D. Supplementary materials with video visualizations of more 3D generations are available at https://brain-3d.github.io/.